A 5.00-kg box slides 3.00 m across the floor before coming to rest. What is the coefficient of kinetic friction between the floor and the box if the box had an initial speed of 3.00 m/s?


1
Expert's answer
2021-10-10T15:56:18-0400

Given:

"m=5.00\:\rm kg"

"d=3.00\:\rm m"

"v_0=3.00\:\rm m/s"


The enegry-work theorem says

"\frac{mv_f^2}{2}-\frac{mv_0^2}{2}=W=-\mu m g d"

So

"\mu=\frac{v_0^2}{2gd}=\frac{3.00^2}{2*9.81*3.00}=0.153"